# Overcoming data challenges to measure whole-person health in electronic health records

### v2

@@ 2.3 Targeted Designs to Select Informative Patients for Validation
As a baseline, consider the simple random sampling (SRS) design, wherein $n$ of the $N$ patients are chosen for validation with equal probability. This strategy ignores all available information, making it the most uninformative. For full-likelihood approaches like the SMLE, any information available for all patients can – and should – be used to design the validation study. Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
@@
When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated log OR on ALI), we start at the end. We begin with the planned logistic regression model of $Y$ given $X$ and $Z$, and work backward to determine which patients will be most informative to it. For this reason, the “best” designs vary across different estimators, with model-based estimators, like the SMLEs, offering the most design flexibility. Other estimators that rely on the design–rather than a model–to handle the missingness, like inverse probability weighting, require nonzero sampling probability for all patients, but model-based estimators do not.
@@
We considered stratified and rank-based validation study designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [32, 26, 33]. Ultimately, simulations based on preliminary validation data (see Section 3.3.3) guided our selection of the targeted design used in the application.
@@
Given our binary outcome, the case-control (CC) design is a natural option here, wherein $n/2$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 2B). The balanced case-control (BCC) design further stratifies on covariate information, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 2C). When the covariates are continuous, they must be discretized to form strata. For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average. Then, $n/4$ patients are validated from each of the four $(Y,X_{D}^{*})$ strata. Importantly, we can still fit the model using the original continuous values.
@@
While the CC and BCC designs are intuitive and can offer some efficiency gains over SRS, there are slightly more complicated designs that are guaranteed to be theoretically “optimal,” meaning that they choose the number of patients to validate from each stratum to achieve the lowest possible variance for the log OR on ALI. Many have been proposed to select patients for the collection of “expensive” binary exposures, like biomarker measurement or validation via chart review, for a variety of analytical approaches to modeling binary outcomes [34, 35, 36, 37, 29, 38]. Recently, Lotspeich et al. (2024) [33] proposed a design for optimal validation of secondary use data for maximum likelihood estimation with binary outcome and exposure misclassification (Figure 2D), which could be applied here. However, we would achieve the lowest possible variance for the discretized version $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., defining more strata) can help, but information will still be lost.
@@
Still, designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains. For example, extreme tail sampling (ETS) designs focus on the extremes (i.e., tails) of the distribution, ranking the data by some value of interest and targeting patients with the smallest or largest values. To identify patients who are most informative not just about ALI ($X^{*}$ and $X$) but about its relationship to healthcare utilization ($Y$), residual sampling (RS) is preferred [39]. With measurement error, we begin by fitting the “naive” analysis model for $Y$ (i.e., using $X^{*}$ instead of $X$) to obtain estimates for the log ORs $\hat{\boldsymbol{\beta}}^{*}$. Then, residuals for the naive model are computed for each patient $i$ ($i=1,\dots,N$) as
@@
$=Y_{i}-\frac{1}{1+\exp\left\{-\left(\beta_{0}^{*}+\beta_{1}^{*}X_{i}^{*}+{\boldsymbol{\beta}}_{2}^{*}Z_{i}\right)\right\}},$ (4)
and used to rank the $N$ patients in the study. With logistic regression, the largest residuals $r_{i}^{*}$ (in magnitude) correspond to observations where the observed value $Y_{i}$ had a low predicted probability (Figure 2E). The auditDesignR package, available on GitHub at www.github.com/sarahlotspeich/auditDesignR, contains R functions to sample from all of these designs.
2.4 Enriched Validation Protocol to Reduce Errors and Missingness
The chart reviews to validate the EHR data were conducted over $6$ months (February to July 2024) by a team of four clinical research assistants and technicians in the Wake Forest University School of Medicine (hereafter, “auditors”). These skilled auditors were chosen for their familiarity with Epic, the institution’s electronic charting software, and their demonstrated experience in the clinical research environment.
@@ 2.4.1 Identifying Sources of Auxiliary Information
Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1). While it is unlikely that we would identify the original missing measurements, we harnessed the clinical expertise of our team to identify sources of auxiliary information about the missing discretized components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values.
For example, if a patient were missing HbA1c, the auditors would search their chart for a diagnosis of diabetes or impaired glycemic control. If either were present, the patient’s diagnosis would suggest their missing HbA1c value would have exceeded the clinically meaningful threshold of $6.5\%$. Thus, their validated ALI $X$ would include one point for HbA1C. If both were absent, the patient’s validated ALI would continue to exclude this still-missing component.
Many components, like C-reactive protein and HbA1c, had multiple sources of auxiliary information, while others, like triglycerides and total cholesterol, had just one. Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
@@ 2.4.2 Validated Data Collection
The EHR data were extracted by the Wake Forest Clinical and Translational Science Institute and securely received by the analytical team. In preparation for validation, the analytical team transformed the data from the original wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable). This transformation streamlined the data structure for the auditors, with only one value to validate per row (Figure 3).
Figure 3: Illustrative example of how the original extracted EHR data for patients’ vitals were transformed from a wide format (with one row per patient encounter and columns for different variables) to a long format (with one row per patient encounter per variable), merged with the audit roadmap, and augmented with empty columns for auditors to enter their findings. The same was done with patients’ lab data.
@@
Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one found in Epic). Depending on the scenario, the auditors would enter different information for the reviewed value, and each data point was assigned an audit finding (flow chart in Figure 4). Three findings were possible for non-missing components in the EHR.
  1. 1.
    Extracted value correct: If the component matched Epic, the auditor would enter the original value again.
  2. 2.
    Extracted value incorrect: If the component did not match Epic, the auditor would enter the new value found in Epic.
  3. 3.
    Extracted value not found: If the component could not be found in Epic, the auditor would enter “Not Found.”
Two findings were possible for missing components.
  1. 4.
    Auxiliary information found: If auxiliary information from the audit roadmap were found in Epic, the auditor would enter the auxiliary information.
  2. 5.
    Auxiliary information not found: If auxiliary information from the audit roadmap were not found in Epic, the auditor would enter “Not Found.”
@@
Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding audit finding assigned by the analysts (Step 4). Column names refer to those in Figure 3.
3 Results
3.1 Cohort Description
We obtained a representative sample of $N=1000$ patients from the EHR at Atrium Health Wake Forest Baptist Hospital in Winston-Salem, North Carolina. To be considered, patients needed to (i) be $18-65$ years old and (ii) initiate their first primary care outpatient encounter in the health system between March 11, 2018, and March 10, 2020 (to avoid data challenges and biases associated with the COVID-19 pandemic). Outpatient encounters were chosen to represent the typical patient voluntarily engaging in care.
Our data contained demographics, vitals, lab values, discharge diagnoses, and problem list values dating $3$ years before the first encounter. Key variables included the measurements for the ten components of the ALI (Table 1), age at first encounter (hereafter, “age”), and an indicator of having at least one hospitalization or ED visit during the $2$-year study period. Since most patients had multiple encounters, the ALI components were based on mean measurements.
Sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, $6\%$ Hispanic, and $73\%$ homeowners. There were $318$ patients ($32\%$) with at least one hospitalization or ED visit. Based on the original unvalidated EHR data, the distribution of error-prone ALI was right-skewed (Supplemental Figure S2). With a median of $0.33$ (interquartile range $=0.17,0.50$), on average, patients in our study experienced $1$ out of every $3$ non-missing stressors on their whole-person health.
3.2 Baseline Data Quality and Naive Model Estimates
As expected, the extracted EHR data had a large amount of missingness. In particular, rare labs for inflammation components were missing for $>95\%$ of patients. Labs for the metabolic components were missing for $21\%$ of patients, and measurements for the cardiovascular components were non-missing for basically everyone (Supplemental Figure S3A). The most common patterns of missing data are displayed in Supplemental Figure S4. Fortunately, most patients ($78\%$) had at least $5$ non-missing ALI components (Supplemental Figure S3B).
Using these ALI values to fit the “naive” healthcare utilization model, we estimate the following. First, for every $0.1$-point (or one component) increase in ALI, the odds of engaging in the healthcare system are expected to increase by $10\%$ (OR $=1.10$, 95% confidence interval [95% CI]: $1.03$, $1.18$), adjusting for age. Second, the expected odds of engaging in the healthcare system for an $18$-year-old patient with an ALI of 00 are $0.25$ (95% CI: $0.18$, $0.35$). Third, for every $10$ years older a patient was at their first encounter, their odds of engaging in the healthcare system are expected to increase by $11\%$ (OR $=1.11$, 95% CI: $1.00,1.23$), adjusting for ALI. Thus, the naive model suggests that worse whole-person health leads to higher expected odds of hospitalization or emergency department visits.
3.3 Three Waves of Data Validation Via Chart Review
While the naive model aligns with the existing literature, we wanted to ensure that the missingness and potential errors in the EHR data did not lead to spurious conclusions. Thus, we secured funding to validate data for $n=100$ of the $N=1000$ patients, whose selection and review were broken down into three waves: (i) Pilot, (ii) Wave I, and (iii) Wave II.
3.3.1 Pilot Wave: Testing the Audit Protocol and Establishing Inter-Auditor Agreement
The first $n_{0}=4$ patients were selected for chart review via the BCC design introduced in Section 2.3. Specifically, patients were split into four strata based on their outcome and error-prone ALI, discretized at the median:
  • (i)
    Engaged in the healthcare system and low error-prone ALI ($Y=1$, $X^{*}\leq 0.33$),
  • (ii)
    Engaged in the healthcare system and high error-prone ALI ($Y=1$, $X^{*}>0.33$),
  • (iii)
    Did not engage in the healthcare system and low error-prone ALI ($Y=0$, $X^{*}\leq 0.33$), and
  • (iv)
    Did not engage in the healthcare system and high error-prone ALI ($Y=0$, $X^{*}>0.33$).
Then, $1$ patient was randomly chosen for validation from each stratum.
The auditors followed the protocol in Section 2.4 to (i) verify the non-missing values in the EHR data or note any discrepancies and (ii) locate auxiliary information for missing values. Following the completion of the Pilot wave, the auditors and analysts met to discuss any modifications that would be helpful before proceeding. For example, the third audit finding of “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded. Adding an optional “Notes” column, where auditors could include any additional information they uncovered, also resulted from this discussion. The Pilot and Wave I findings were combined to interpret data quality (see Section 3.3.2).
We started with two auditors on our team. Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined the team, each reviewing these four patients’ records as well. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [40], as detailed in the Supplemental Materials. We found high agreement between the auditors’ findings for lab ($\kappa=0.90$) and vitals ($\kappa=0.81$) variables.
3.3.2 Wave I: Collecting Preliminary Audit Data
@@
The next $n_{1}=48$ patients validated were also selected via the BCC design from Section 3.3.1. Since we did not have any prior information about the quality of the EHR data, we sampled more conservatively here to collect preliminary information for a more extreme, targeted design in Wave II. Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [32, 26]. Even if prior information is available (e.g., from a previous audit), validating in multiple waves allows us to relearn how our data may have changed before using all our resources to validate them [33].
[…]
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable.
Fewer vitals than labs were found to match Epic ($59\%$ versus $81\%$), with $31\%$ not found. Through validation, we discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. For encounters where vitals were not taken (e.g., physical therapy or optometry visits), values were imputed from other encounters; these imputed values could not be found in Epic by the auditors. Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
Incorporating the audit roadmap into the validation protocol improved the completeness of the data, as the number of missing values decreased for nearly all components of the ALI (Supplemental Figure S3A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of the $52$ validated patients ($44\%$). The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average. Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
Incorporating the Pilot and Wave I, we re-fit the healthcare utilization model using the SMLE with the first $52$ validated and $948$ unvalidated patients’ data (Table 2). The association between whole-person health and healthcare utilization appeared even stronger, with an estimated 21% increase in odds for each $0.1$-point increase in ALI (OR $=1.21$, 95% CI: $1.17,1.25$), adjusting for age. For an 18-year-old patient with an ALI of 00, the expected odds of engaging in the healthcare system decreased to $0.17$ (95% CI: $0.12$, $0.26$). The adjusted OR for age was relatively unchanged (OR $=1.10$, 95% CI: $1.01,1.21$). Thus, the preliminary SMLE model reaffirms that worse whole-person health is associated with higher expected odds of engaging in the healthcare system.
Odds/Odds
Coefficient Ratio Estimate 95% CI
Naive Analysis (Unvalidated Data)
Intercept $0.251$ $(0.176,0.357)$
ALI $1.099$ $(1.028,1.175)$
Age $1.108$ $(1.000,1.228)$
Preliminary SMLE (Pilot + Wave I Validated Data)
Intercept $0.173$ $(0.115,0.262)$
ALI $1.207$ $(1.167,1.247)$
Age $1.104$ $(1.007,1.211)$
Final SMLE (All Waves Validated Data)
Intercept $0.225$ $(0.156,0.327)$
ALI $1.120$ $(1.088,1.154)$
Age $1.105$ $(1.018,1.200)$
Table 2: Estimates and 95% confidence intervals for coefficients from the naive analysis (logistic regression, no validated data), preliminary sieve maximum likelihood estimator (SMLE, validated data on $52$ patients), and final SMLE (validated data on $100$ patients) models.
3.3.3 Simulations: Using Preliminary Audit Findings to Design Wave II
@@
Before validating the remaining patients, we designed bespoke simulation studies based on the preliminary audit data. These simulations allowed us to explore how potential designs compared under different scenarios for the true underlying data quality. In particular, the design offering the best statistical efficiency for the OR on ALI was desired. Focusing here on the takeaways and our decision-making, details like the data-generating mechanisms can be found in the Supplemental Materials. The five designs described in Section 2.3 were considered: SRS, CC, BCC, OPT, and RS.
Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency (i.e., smaller variance) than the uninformative SRS strategy. The RS design provided the best efficiency under extra low to moderate error settings; the preliminary validation data indicated that we had a perfect TPR $=1.00$ and negligible FPR $<0.01$.
@@
The SMLE $\hat{\beta}_{1}$ was empirically unbiased under full/high data recovery rates ($90-100\%$ missing data recovered through validation) and different validation study designs. While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on the unvalidated EHR data). The RS design offered the highest efficiency, regardless of how much missing data was recovered. Thus, based on extensive simulations considering different error rates and data recovery scenarios, we concluded that the RS design was the most informative way to select the remaining $n_{2}=48$ patients for validation in Wave II.
3.3.4 Wave II: Implementing the Targeted Validation Study Design
@@
For the RS design, each unvalidated patient’s residual was calculated based on the naive model estimates, following (4), as

### v4

@@ 2.4 Design: Targeted Sampling to Select Informative Patients for Validation
Based on our analysis plan, we considered various sampling designs to select the most informative $100$ patients for validation. When validating computable phenotypes, simple random sampling (SRS) is often used. Under SRS, $100$ patients would be chosen for chart review with equal probability (Figure 3A). This strategy ignores all available information. Some phenotyping algorithms require an SRS validation study. However, for full-likelihood approaches like SMLEs, any information available for all patients can be used to design the validation study. Designing based on fully observed information ensures that the MAR assumption holds for the downstream analysis while allowing for many targeted strategies.
Figure 3: Diagram of partial validation studies with $N$ patients extracted from the electronic health records (EHR) and $n$ patients chosen for validation via chart review. The least informative strategy to select the validation study is A) simple random sampling. More informative targeted strategies include B) case-control sampling, C) balanced case-control sampling using discretized ALI, D) optimal stratified sampling using discretized ALI, and E) residual sampling based on the naive model. Within strata, the darker shaded areas indicate which patients would be selected.
@@
When designing a validation study to promote statistical efficiency of the parameter estimates (i.e., to minimize the variance in the estimated logOR on ALI), we start at the end. Beginning with the planned logistic regression model of $Y$ given ($X$, $Z$), we work backward to determine which patients will be most informative. For this reason, the “best” designs vary across different estimators, with model-based ones, like SMLEs, offering the most design flexibility. Estimators that rely on the design to handle missingness, like inverse probability weighting (IPW), require nonzero sampling probability for all patients; model-based estimators do not.
@@
We considered stratified and rank-based designs. Multiple validation waves were conducted to allow the protocol and study design to adapt as information accumulated, which is becoming common in this space [42, 34, 43]. Ultimately, simulations based on preliminary validation data guided our selection of the targeted design used in the application (see Section 2.6).
2.4.1 Stratified Designs
@@
Given our binary outcome, the case-control (CC) design is a natural option, wherein $50$ patients each are chosen from subgroups with $Y=0$ and $Y=1$ (Figure 3B). The balanced case-control (BCC) design further stratifies on covariates, subsequently choosing equal (i.e., balanced) numbers from each stratum (Figure 3C). Continuous covariates must be discretized to form strata. For instance, we could use $X_{D}^{*}$, a discretized version of $X^{*}$, to define strata of patients with error-prone ALI above and below the average. Then, validate $25$ patients from each of the four ($Y$, $X_{D}^{*}$) strata. Importantly, we can still fit the model using the continuous values $X^{*}$.
@@
The CC and BCC designs are intuitive and offer efficiency gains over SRS. Still, slightly more complicated designs are guaranteed to be theoretically “optimal;” they choose the number of patients from each stratum to achieve the lowest possible variance for the logOR on ALI [44, 45, 46, 47, 37, 48]. [43] proposed the optimal (OPT) design for model-based analysis of data with binary outcome and exposure misclassification (Figure 3D), which could apply here. However, we would achieve the lowest possible variance for $X_{D}^{*}$ rather than $X^{*}$. Increasing the number of unique values of $X_{D}^{*}$ (i.e., more strata) can help, but information will still be lost. Moreover, the OPT design depends on the true logORs, which are unknown; reliable placeholders for them are required to implement the design (often collected in a multi-wave framework).
2.4.2 Rank-Based Designs
@@
Designs that rank patients rather than stratifying them offer improved granularity and sizable efficiency gains. For example, extreme tail sampling (ETS) designs rank patients by some value of interest and target those with the smallest or largest (i.e., most “extreme”) values. To identify patients most informative not just about ALI but its relationship to healthcare utilization, residual sampling (RS) is preferred [49]. To implement RS with measurement error, we begin by fitting the “naive” analysis model, i.e., using standard logistic regression procedures with $Y$, $X^{*}$ (in place of $X$), and $Z$ to obtain estimates for the logORs $\hat{\boldsymbol{\beta}}^{*}$ from the full sample of $1000$ patients. Then, the naive model residuals are computed for each patient $i$ as
@@
$=Y_{i}-\left[1+\exp\left\{-\left(\hat{\beta}_{0}^{*}+\hat{\beta}_{1}^{*}X_{i}^{*}+\hat{\beta}_{2}^{*}Z_{i}\right)\right\}\right]^{-1},$ (4)
and used to rank patients. The largest residuals $r_{i}^{*}$ (in magnitude) correspond to patients whose observed outcome $Y_{i}$ had a low predicted probability (Figure 3E). The auditDesignR package contains R functions to sample from all of these designs [50].
2.5 Protocol: Enriched Validation Protocol to Reduce Errors and Missingness
Chart reviews were conducted over $6$ months by four clinical research assistants and technicians at WFUSM (hereafter, “auditors”), chosen for their familiarity with Epic (the institution’s electronic charting software) and demonstrated research experience.
@@ 2.5.1 Identifying Sources of Auxiliary Information
Before the chart reviews, our clinical collaborators developed the “audit roadmap” (Table 1). While it is unlikely that we would identify the original missing measurements, we harnessed clinical expertise to identify auxiliary information about missing discretized components. The roadmap suggested where auditors could look in Epic to locate supplemental information about missing values. It employs “anchors” to try and recover missing components using clinically-driven placeholders [15].
For example, if a patient were missing HbA1c, auditors would search their chart for a diagnosis of diabetes or impaired glycemic control. If either was present, the patient’s diagnosis would suggest that their missing HbA1c value would have exceeded the $6.5\%$ threshold. Thus, their validated ALI $X$ would include one point for HbA1c. If neither diagnosis was present, $X$ would continue to exclude this still-missing component.
Many components, like HbA1c, had multiple sources of auxiliary information, while others, like triglycerides, had just one. Only serum albumin, a blood test to measure liver and kidney function, did not have a supplemental diagnosis.
@@ 2.5.2 Validated Data Collection
In preparation for validation, the analytical team transformed the extracted EHR data from the original wide format (one row per patient encounter and columns for different variables) to a long format (one row per patient encounter per variable). This transformation streamlined the data structure for auditors: one value to review per row (Supplemental Figure S2).
@@
Before sending the data to the auditors, the roadmap was incorporated, and an empty column was added for the “reviewed value” (i.e., the one in Epic). Depending on the scenario, auditors would enter different information for the reviewed value, and each data point was assigned one of five auditor findings (Figure 4).
@@
Figure 4: Flow chart of instructions given to auditors for the chart reviews (Steps 1–3) and the corresponding auditor finding assigned by the analysts (Step 4). Column names refer to those in Supplemental Figure S2.
2.6 Multi-Wave Strategy: Sampling Adaptively and Improving Tools
@@
Choosing the “best” design often relies on information that is unknown at baseline, like EHR error rates or the magnitude of the association between ALI and healthcare utilization. Since we did not have any prior information about these unknowns, we started by sampling conservatively to collect preliminary information for a more extreme, targeted design later on. Multi-wave validation schemes like this one are appealing because they balance learning about our data with strategically designing for them [42, 34]. Even if prior information is available (e.g., from previous chart reviews), multi-wave validation allows us to relearn how data have changed before dedicating resources to validate them [43]. Further, we can seize opportunities to incorporate auditors’ feedback and improve the validation protocol.
We validated $100$ patients’ data across three waves: the Pilot, Wave I, and Wave II. The small Pilot wave tested the validation protocol and established inter-auditor agreement on the first four patients. Then, Waves I and II put this protocol into action, as auditors validated the remaining patients. Waves I and II each selected $48$ more patients. (Designating equal-sized waves balances design flexibility and stability [42].) In Wave II, information from the Pilot + Wave I was leveraged to design a more targeted strategy. Patients could be chosen for validation at most once.
2.6.1 Adapting the Validation Study Design as Information Accumulates
Between the Pilot and Wave I, the first $52$ patients were selected for chart review via the BCC design. Specifically, patients were split into four strata based on their outcome and error-prone ALI (discretized at the median of $0.33$). Then, equal numbers of patients were randomly chosen from the four strata. In the Pilot, one patient was chosen from each stratum; another $12$ were chosen per-stratum in Wave I.
@@ 2.6.1 Adapting the Validation Study Design as Information Accumulates
Before validating the remaining $48$ patients, we designed bespoke simulation studies based on the preliminary validation data. They explored how designs compared under different scenarios for true underlying data quality. The design offering the best efficiency (i.e., lowest variability) in estimating the logOR on ALI with the SMLE was desired.
@@
Under varied error rates, the targeted validation study designs offered empirically unbiased estimates and higher efficiency than SRS. The RS design provided the best efficiency under extra low to moderate error settings, and the preliminary validation data (Pilot + Wave I) indicated a perfect true positive rate (TPR) $=1.00$ and negligible false positive rate (FPR) $<0.01$. Under full/high data recovery rates ($90$-$100\%$ missing data recovered through validation) and different designs, the SMLE was empirically unbiased. While biased for moderate recovery or lower ($\leq 50\%$), the SMLE was still closer to the truth than the naive estimate was (i.e., based on unvalidated EHR data). Regardless of how much missing data was recovered, the RS design offered the highest efficiency. Details like the data-generating mechanisms can be found in the Supplemental Materials.
@@
We concluded that the RS design was the most informative way to select the remaining $48$ patients for validation in Wave II. To implement it, each unvalidated patient’s residual was calculated based on the naive healthcare utilization model estimates (Section 3.3) following (4), as
@@
2.6.2 Testing the Validation Protocol and Establishing Inter-Auditor Agreement
The validation protocol evolved slightly throughout the waves, as well. After completing the first four patients’ chart reviews (Pilot), auditors and analysts met to discuss any protocol modifications that would be helpful before proceeding. For example, the auditor finding “extracted value not found” was not one of the original options, and the team discussed how these data points should be recorded. Adding an optional “Notes” column, where auditors could include any additional information, also resulted from this discussion.
The Pilot also presented a valuable opportunity to evaluate inter-auditor agreement. We started with two auditors on our team. Each reviewed the four patients chosen for the Pilot, allowing us to calculate inter-auditor agreement in this doubly-validated sample. Later, two new auditors joined the team, each reviewing these four patients’ records as well. Ultimately, agreement in the four auditors’ findings was calculated using Fleiss’ Kappa [51], as detailed in the Supplemental Materials. We found high agreement between the auditors’ findings for labs ($\kappa=0.90$) and vitals ($\kappa=0.81$).
3 Results
3.1 Cohort Description and Study Objectives
We considered adult patients ($18$-$65$ years old) with at least one primary care outpatient encounter at Atrium Health Wake Forest Baptist Hospital. The $1000$ sampled patients had a median age of $48$ years and a median income of \$52,320. They were $61\%$ female, $72\%$ White, and $6\%$ Hispanic. During the $2$-year study period, $318$ patients ($32\%$) had at least one hospitalization or ED visit. In unvalidated EHR data, error-prone ALI was right-skewed (Supplemental Figure S3). With a median of $0.33$ (interquartile range $=0.17,0.50$), patients experienced one out of every three non-missing components on their whole-person health, on average. We validated ALI data for $100$ patients.
First, we discuss the EHR data quality (Section 3.2). Second, we present the healthcare utilization model (Section 3.3). Within these sections, results are discussed chronologically, beginning with findings from the unvalidated EHR data and ending with those following the completion of all validation waves. We also discuss how these findings inform future implementations of the ALI in this healthcare system (Section 3.4).
3.2 Validation Study Findings: EHR Data Quality
As expected, the extracted EHR data had lots of missingness. Rare labs for inflammation components were missing for $>95\%$ of patients (Supplemental Figure S4A). Labs for metabolic components were missing for $21\%$ of patients. Common patterns of missing data are displayed in Supplemental Figure S5. Measurements for the cardiovascular components were non-missing for basically everyone. Fortunately, most patients ($78\%$) had at least five non-missing ALI components (Supplemental Figure S4B).
3.2.1 Pilot + Wave I Chart Reviews
[…]
Figure 5: Auditors’ findings from the validation of A) $52$ patients’ data in the Pilot + Wave I, B) $48$ patients’ data in Wave II, and C) $100$ patients’ data in all waves combined. These findings refer to the original numeric measurements (before discretizing them into allostatic load index components), and there could be multiple per patient per variable. Auditor findings were mutually exclusive and specific to whether the measurements were originally non-missing (“extracted value correct,” “extracted value incorrect,” or “extracted value not found”) or missing (“no auxiliary information found” or “auxiliary information found”) in the extracted EHR data.
Fewer vitals than labs matched Epic ($59\%$ versus $81\%$), with $31\%$ not found. We discovered that the EHR data extraction algorithm expected every patient encounter to have vitals. For encounters where vitals were not taken (e.g., physical therapy), values were imputed from other encounters; imputed values could not be found in Epic by the auditors. Thus, we uncovered an area for potential improvement in the current EHR data extraction protocol.
Incorporating the audit roadmap into the validation protocol improved the completeness of the data. The number of missing values decreased for nearly all ALI components (Supplemental Figure S4A). Homocysteine saw the biggest improvement, with auxiliary information located for $23$ of $52$ validated patients ($44\%$). The number of non-missing ALI components per patient increased from $6.6$ to $7.5$, on average. Overall, $48$ of $177$ missing components ($27\%$) were recovered through validation (Figure 6A).
3.2.2 Wave II Chart Reviews